OpenStudy (anonymous):

THE SUM OF THE INTERIOR ANGLES OF THE POLYGON IS 2520 DEGREE FIND THE NUMBER OF SIDES

OpenStudy (anonymous):

2520=(n-2)180

OpenStudy (anonymous):

It's easy see.... (n-2)(180)=2520 180n-360=2520 add 360 to both sides 180n=2880 n=16 sides

OpenStudy (anonymous):

Hope it helps :O

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

yeah np, Anytime buddy )

OpenStudy (anonymous):

That's the formula u should remember........ Sum of Interior Angles of a Polygon= 180(n - 2) :D